1. Indicate which of the following are..

i. Choose the degree of elasticity from the menu that matches the statement: Income changes from 80,000 to 96,000 and consumption of an item changes from 50 to 45 Options Unit Elastic, Perfectly Elastic, Elastic, and Inelastic

ii. Choose the degree of elasticity from the menu that matches the statement: The price of T-bone steaks changes from 4/lb to 4.8/lb, the demand of sirloin steak changes from 50 lbs to 55 lbs Options: Inelastic, Elastic, Unit Elastic, Perfectly Elastic

iii. Choose the degree of elasticity from the menu that matches the statement: When the price of lemons changed from 0.85 to 1.02, the demand for tea went from 20 to 18 units Options: Perfectly Elastic, Unit Elastic, Inelastic, Elastic
Answer #1

i) E = [45-50/(45+50/2)]/[96000-80000/(96000+80000/2)]

= [-5/47.5]/[16000/88000]

= 0.1053/0.1818

= 0.58

Inelastic

ii) E = [55-50/(55+50/2)]/[4.8-4/(4.8+4/2)]

= [5/52.5]/[0.8/4.4]

= 0.0952/0.1818

= 0.52

Inelastic

iii) E = [18-20/(28+20/2)]/[1.02-0.85/(1.02+0.85/2)]

= [2/24]/[0.17/0.935]

= 0.0833/0.1818

= 0.46

Inelastic